### v1

@@ 3.7.2 Detection units settings
To configure our detection units, we must first define our target boxes. The True catalog does not contain the necessary information to define them in the classical computer vision way. Instead, we define our boxes as centered on the source central coordinates and with a size that is scaled on its major and minor axes. For each target source, we first define a rectangular box with $\hat{w}_{PA=0}=2\times Bmaj$ and $\hat{h}_{PA=0}=2\times Bmin$ corresponding to $PA=0$. This box is then rotated to correspond to the actual $\hat{PA}$ of the source, and we search for the smallest square box that contains the four rotated vertex. The resulting dimensions are clipped in the 5 to 64 pixels range to obtain the final $\hat{w}$ and $\hat{h}$ dimensions that are used to define our target box for the corresponding source (Fig. 9). The minimum clipping implies that all unresolved point sources get the same minimum size. We stress that in our specific context, box sizes do not have to be very accurate. Firstly, they do not constrain the receptive field in any way, meaning that the network can use information outside the box to make the detection. Secondly, they are only used during association so the detector can be trained, the important $Bmaj$ and $Bmin$ values being predicted as additional parameters (Sect. 3.7.3).
@@
We define the number of detection units and their size priors based on the target source density and their box-size distribution. We chose to have three size regimes: i) a small regime that is composed of several detection units with an identical $6{\times}6$ size prior, ii) an intermediate regime with two detection units with the same surface but two aspect ratios with size priors of $9{\times}12$ and $12{\times}9$ respectively, and iii) a large regime to represent all sources up to our maximum size clipping value with a single detection unit with a $24{\times}24$ size prior. We illustrate how the target sources distribute over these size-priors based on the smallest Euclidian distance with their respective size Fig. 12. It indicates that most of the sources would be theoretically associated with the smallest size regime. From this, we tried having all our detection units in the small regime, which resulted in bad detection accuracy for all target sizes. This confirms that the source size remains an appropriate first-order criterion for distributing the network expressivity over the detection units, even with such a massive target size regime imbalance.
@@
With the small-source density being high, we need to be capable of detecting several of them in the same grid cell. For this, we use several identical detection units in the small regime. We searched for the optimum number of $6{\times}6$ detection units and observed good results in the 4 to 8 range. Too few detection units limit the number of detectable objects, and it might force them to encapsulate multiple contexts, preventing them from being detected simultaneously at prediction time. Conversely, too many boxes dilute the context diversity and increase the training difficulty. We settled our detector to have 6 small detection units. With the 3 detection units from the two other size regimes, our detector predicts a total number of 9 independent boxes in the same grid cell. We will discuss the fact that all these detection units are never used simultaneously in practice and how the actual predictions are distributed among them in Sect. 4.2.
@@
Figure 12: Target distribution over the closest box size priors. The association is made using the Euclidean distance in the 2D size space.
@@ 3.7.3 Source properties to predict
With our method, regardless of the application, each detection unit must always predict the box center coordinates ($x,y$), the box size ($w,h$), and one of or both the probability $P$ and the Objectness $O$. For all the remaining parameters requested by the SDC1 task, namely the source flux $f$, the major axis size $Bmaj$, the minor axis size $Bmin$, and the position angle $PA$, we can use the extra-parameter prediction capability of YOLO-CIANNA (Sect. 2.4). As stated in Sect. 3.1, we choose to ignore source classification and the associated core fraction and, therefore, do not predict these properties. Still, the scorer requires values for these parameters to compute the per-source score. We then adopt the simple approach of setting a constant value for these parameters for all sources. We found that setting $cf=0.0375$ and $C=2$ (corresponding to SFG) resulted in the best average sub-scores over the full catalog with $\bar{s}^{cf}=0.9865$ and $\bar{s}^{C}=0.9734$.
@@
For all the extra parameters, their output activation is linear (Fig. 3), and we have to normalize them to be all in a similar range. For the source flux, we apply minimum and maximum clipping limits so $f$ is in the $[1.9{\times}10^{-6},2{\times}10^{-3}]$ Jy range, followed by a log10 rescale. The resulting distribution is then linearly rescaled in the 0 to 1 range using the clipping limits as interval edges. We stress that our target is the apparent flux in Jy, obtained by multiplying the integrated source fluxes from the True catalog by the interpolated primary beam value at the source central coordinates. $Bmaj$ and $Bmin$ follow the same normalization scheme, with $Bmaj$ in the $[0.9,60.0]$ arcsec range and $Bmin$ in the $[0.3,30.0]$ arcsec range. For $PA$, it is very important to consider if the source is resolved or not. Trying to predict the position angle of an unresolved source would result in training noise. We first search for all sources with $Bmaj\leq 1.8$ arcsec and attribute them a target of $\hat{PA}=0$. We note that this filtering only applies when defining the target and that nothing prevents the detector from predicting an angle for smaller sources at prediction time. We then define two predicted parameters, $\cos{(PA)}$ and $\sin{(PA)}$. Considering that the target $PA$ is in the $[-90,90]$ degree range, we linearly rescale $\sin{(PA)}$ to the 0 to 1 range. Due to angular symmetries and degeneracies of PA, predicting both $\cos{(PA)}$ and $\sin{(PA)}$ with the network and then reconstructing the predicted $PA$ to add to the source catalog works better than a direct angle prediction. In summary, each box predicts the following vector for each source $x,y,w,h,P,O,f,Bmaj,Bmin,\cos{(PA)},\sin{(PA)}\right>$. As discussed in Sect. 2.4, we can set independent $\gamma^{p}$ scaling values for all parameters to balance their respective importance in the loss. We list the $\gamma^{p}$ we used for all parameters in Table 3.
3.7.4 Remaining hyper-parameters
@@
We list all the hyper-parameters for our application here for reproducibility purposes, even if some of them are only described in the appendix Appendix A.3, with the advanced description of the association process of YOLO-CIANNA. We use the alternative ${\rm DIoU}$ matching metric, described in Appendix. A.4, which implies that the $L^{{\rm fIoU}}$ limits can take values in the -1 to 1 range. Following our nine detection units definition with 3 scales, we set $\lambda_{void}=0.15$ for the 6 small detection units and $\lambda_{void}=0.01$ for the 3 larger ones. The different $\lambda$ scaling factors, along with the corresponding pre-activation scaling and limit values, are given in Table 2. The various $L^{{\rm fIoU}}$ and some other remaining hyperparameters are given in Table 3. For a more exhaustive view of all the detector hyperparameters, we recommend reading the provided example SDC1 scripts archived with CIANNA V1.0.
Table 2: List of scaling and limits related to the loss sub-parts.
@@
Table 3: List of some of the detector hyperparameters.
Parameter $\gamma^{f}$ $\gamma^{Bmaj}$ $\gamma^{Bmin}$ $\gamma^{cos{(PA)}}$ $\gamma^{\sin{(PA)}}$
@@
Parameter $L^{{\rm fIoU}}_{gbnb}$ $L^{{\rm fIoU}}_{low}$ $L^{{\rm fIoU}}_{P\_good}$ $L^{{\rm fIoU}}_{O\_good}$ $L^{{\rm fIoU}}_{p\_good}$
@@
Parameter $S_{ar}$ $\alpha_{small}$ $N_{rand}$ $\alpha_{best}$ $\alpha_{rand}$
@@
3.8 Detector training
@@
The layer weights are initialized to random values following a Glorot normal initialization (Glorot & Bengio, 2010), the batch size is set to $b_{s}=16$, and the numerical resolution is set to full 32-bit floating point for all network elements. Our default learning rate is $1.5{\times}10^{-4}$ with an exponential decay of $5{\times}10^{-4}$ as a function of the training iteration up to a minimum value of $5\%$ of the default learning rate. A constant weight update momentum of 0.8 is also used. No weight decay is used since it always resulted in lower scores. The first 40 iterations are done using an increasing learning rate up to the reference value, and the initial random association is set to $N_{rand}=16000$ (Appendix A.6.2).
@@
With our fully convolutional architecture, we could technically use any input size that is a multiple of the network reduction factor (see Appendix A.1). A too-small input size can cause many cutout edge effects, and a too-large input size reduces the number of examples that can be constructed from a finite training area, increasing the chances of overtraining. For our specific application, we observed that a training input size of $256{\times}256$ pixels resulted in the best achievable score. During training, we dynamically generate groups of 1600 augmented images that we refer to as a training iteration. For each example, we draw a random pixel position inside the training area, extract a small region corresponding to our network input size, and apply multiple pixel-preserving operations randomly from vertical or horizontal flips and $90$ or $-90$ degree rotations. Due to how astronomical images are constructed, any transformation that resamples the pixels must be avoided. Images from the additional noise region are selected at a low $5\%$ rate and follow the same augmentation rules. By cutting out patches of the larger image, sources on the edges of a patch might be cut, which can have a strong negative impact on the detector capabilities (see Sect. 3.4). We tried various approaches for handling these sources (Appendix A.6.4) but obtained the best scores with a simple exclusion of any target box that is not fully contained in the cutout.
@@
Because the labeled area is small, we need to preserve all of it in our training sample. Our validation dataset is then defined from the same region and catalog but using a fixed grid of non-overlapping patches for a total of 100 images. We acknowledge that this is non-optimal as it would fail to identify overtraining. However, we can rely on the SDC1 scorer to provide an independent metric on a separate dataset to evaluate our detector performances at a regular iteration interval. We also use the score over the full image field to optimize the individual objectness limits for all detection units from a trained network. We emphasize that this optimization is a simple refinement and that we would still achieve more than $95\%$ of our best score with naive guesses on these values instead of searching for the best ones.
@@
With this setup, the detector reaches a high score after only 400 iterations, but the larger sources are not properly constrained, and the global characterization is sub-optimal. The detector usually requires around 3000 iterations to converge. After that, it oscillates for a few hundred iterations and eventually exhibits overtraining, lowering the score by up to $5\%$. We present the validation loss with scoring stamps every 200 iterations in Fig. 13. We note that using stricter values for parameters that control the association process like $L^{{\rm fIoU}}_{gbnb}$ or $L^{{\rm fIoU}}_{low}$ would speed up the training. It would result in more stable results, but it usually decreases the best achievable score by a small amount.
@@
On an RTX 4090 GPU and by doing image augmentation on the CPU in parallel, we can reach almost 400 images per second of training performances. Reaching iteration 3000 takes around 3.5 hours. With the current training setup, the GPU memory footprint of the network is around 8GB. A save file representing the model at a given iteration requires 50MB of storage. We tried using different types of mixed-precision along with Nvidia AI-dedicated tensor-cores acceleration for training, but it always resulted in a decrease in score of about $2\%$. Still, it speeds up network training compute performance by a factor of up to $1.8$ while preserving the relative performance impact of various hyperparameters, allowing a more efficient exploration of network architectures and hyperparameter combinations.
@@ 3.9 Prediction pipeline
Table 4: First NMS rejection limit pairs.
$L^{{\rm fIoU}}_{NMS}$ 0.05 -0.1 -0.3 -0.5
$L^{Obj}_{NMS}$ 1.0 0.7 0.5 0.3
@@
To use our trained detector on the full SDC1 image, we need to decompose it into patches. As before, due to the fully convolutional structure of our detector and its resolution static prediction (Appendix A.1), we can technically use any input size for prediction. As for training, decomposing the image into patches results in sources being close to the edges of a patch, which are then poorly detected or characterized. To overcome this at prediction time, we use overlapping patches with an overlap size of 32 pixels, corresponding to the size of two grid cells and half our maximum clipping size for our target boxes. Each source is then more likely to be well represented by at least one patch. In case of multiple detections of the same source by two patches, the objectness score can be used to select the best one through an inter-patch NMS process. From there, we observed that larger patches tend to produce better results as they reduce the proportion of sources close to an edge. We settled for a prediction input size of $512{\times}512$ pixels, which results in a good balance between prediction quality and compute efficiency. By setting a small offset on the edges of the full image, we can obtain a map made of $67{\times}67$ partially overlapping patches.
@@
For each of them, we first apply a per-detection-unit objectness threshold filtering. The threshold values are obtained through score optimization by identifying the objectness interval contributing positively to the score (Sect. 3.3 and Eq. 22). We also add a rejection criterion based on the average pixel flux value in the box area as a function of the predicted objectness score to exclude many small false detections induced by very large and bright sources. Multiple detections are then filtered by the NMS process for each patch (Sect. 2.8), using multiple $L^{{\rm fIoU}}_{NMS}$ and $L^{Obj}_{NMS}$ thresholds pairs that are given in Table 4. Following the NMS process, considering the best current score box, any other box that respects one of the condition pairs ${\rm DIoU}>L^{{\rm fIoU}}_{NMS}$ and $O<L^{Obj}_{NMS}$ is removed. With this setup, the rejection chance is based on a combination of the detection confidence and the distance to other confident detection. To filter multiple detections from overlapping patches, we use a secondary inter-patch NMS process. It works exactly as the first NMS but with a constant $L^{{\rm fIoU}}_{NMS}=-0.15$. From the filtered box list, we can construct a catalog in the scorer format by inverting our normalizations on the predicted parameters and also convert the predicted apparent flux back to a real flux using the interpolated primary beam value at the position of the detection.
Figure 13: Evolution of the validation loss sub-parts during training, using the ”natural” loss representation (Appendix A.7). The first 20 iterations are skipped. Optimized scores obtained using the procedure described in Sect. 3.9 are displayed every 200 iterations (one or two points differences can be observed compared to Table 5 due to rounding of the objectness thresholds).
@@
On an RTX 4090 GPU, the raw prediction compute performance at the largest available batch size is around 300 images per second for an input size of $512{\times}512$ pixels, or almost 80 million pixels per second. The raw processing time for the full 4GB image of 32768 square pixels after the network loading is around 15 seconds. To this, we need to add the post-processing time of the prediction pipeline, which strongly depends on the minimum objectness threshold values for the different detection units. With an end-to-end Python processing pipeline that only produces the converted detection catalog, this processing time can be reduced to approximately 40s on a single CPU core (Ryzen 9 5900X). Interestingly, the detection performances are fully preserved when doing the prediction in mixed-precision using 16-bit numerical computations with a 32-bit accumulator when necessary, allowing the use of Nvidia AI-dedicated tensor-cores compute units available on modern GPU architectures. With this, we can reach around 500 images of size $512{\times}512$ per second, or around 130 million pixels per second, of raw computing performance. At this point, saving the results on our local high-end SSD storage becomes the limiting factor, and the network loading time becomes comparable to the prediction time over the full image. This confirms that our method is perfectly suited for very large surveys and real-time source detection (Sect. 5.3).

### v2

@@ 3.7.2 Detection units settings
To configure our detection units, we must first define our target boxes. The True catalog does not contain the necessary information to define them in the classical computer vision way. Instead, we define our boxes as centered on the source central coordinates and with a size that is scaled on its major and minor axes. For each target source, we define a rectangular box with $\hat{w}_{\textrm{PA}=0}=2\times\textrm{Bmaj}$ and $\hat{h}_{\textrm{PA}=0}=2\times\textrm{Bmin}$ corresponding to $\textrm{PA}=0$. This box is then rotated to correspond to the actual $\hat{\textrm{PA}}$ of the source, and we search for the smallest square box that contains the four rotated vertex. The resulting dimensions are clipped in the 5 to 64 pixels range to obtain the final $\hat{w}$ and $\hat{h}$ dimensions that are used to define our target box for the corresponding source (Fig. 9). The minimum clipping implies that all unresolved point sources get the same minimum size, which was calibrated to optimize the association process of our detection layer. We stress that box sizes do not have to be very accurate in our specific context. Firstly, they do not constrain the receptive field in any way, meaning that the detector can use information outside the box to detect or characterize the source. Secondly, they are only used during association so the detector can be trained and are not used in the scorer matching criteria and characterization score. The important Bmaj and Bmin values are instead predicted as additional parameters (Sect. 3.7.3).
@@
We define the number of detection units and their size priors based on the target source density and the box-size distribution. We chose to have three size regimes: i) a small regime that is composed of several identical units of $6{\times}6$ size prior, ii) an intermediate regime with two units of identical surfaces but two aspect ratios with size priors of $9{\times}12$ and $12{\times}9$ respectively, and iii) a large regime with a single unit of $24{\times}24$ size prior. We illustrate how the target sources distribute over these size-priors based on the smallest Euclidian distance with their respective size in Fig. 12. This indicates that most sources would theoretically be associated with the smallest size regime. We tried an alternative setup with all our detection units in the small regime, but it resulted in lower detection performances for all size regimes. This confirms that the source size remains an appropriate first-order criterion for distributing the network expressivity over the detection units, even with such a massive target size regime imbalance (Appendix A.3).
@@
To detect multiple small sources in the same grid element, we must populate the small-size regime with several identical detection units. We observed good results for all models trained using four to height small units with optimum results for six. Too few detection units limit the number of detectable objects. It can also force some units to encapsulate multiple contexts, preventing them from being detected simultaneously at prediction time. Conversely, too many boxes dilute the context diversity and increase the training difficulty. Combined with the three detection units from the two other size regimes, our detector predicts nine independent boxes in the same grid cell. In practice, all these detection units are never used simultaneously. We discuss how the actual predictions are distributed among them in Sect. 4.2.
@@
Figure 12: Target distribution over the closest box size prior. The association uses the Euclidean distance in the 2D box size space.
@@ 3.7.3 Source properties to predict
With our method, source characterization is handled directly by the detector in the form of a single-stage network. We discuss the possible impact of this design choice in Sect. 5.1.4 and Appendix D. In practice, each detection unit must always predict the box center coordinates ($x,y$), the box size ($w,h$), the probability $P$, and the Objectness $O$. For all the remaining parameters requested by the SDC1 task, namely source flux $f$, major axis size Bmaj, minor axis size Bmin, and position angle PA, we can use the extra-parameter prediction capability of YOLO-CIANNA (Sect. 2.4). As stated in Sect. 3.1, we ignore source classification and the associated core fraction and do not predict these properties. Still, the scorer requires values for these parameters to compute the per-source score. We adopted the simple approach of setting a constant value for these parameters for all sources. We found that setting $cf=0.0375$ and $C=2$ (corresponding to SFG) resulted in the best average subscores over the full catalog with $\bar{s}^{cf}=0.9865$ and $\bar{s}^{C}=0.9734$.
@@
The output activation for all extra parameters is linear (Fig. 3), so we have to normalize in a similar range. For the flux, we convert it to apparent flux as described in Sect. 3.2. Then we apply minimum and maximum clipping limits so $f_{a}$ is in the $[1.9{\times}10^{-6},2{\times}10^{-3}]$ Jy range, which is then passed through a log10 function. The resulting distribution is then linearly rescaled in the 0 to 1 range using the clipping limits as interval edges. For scoring, the predicted apparent flux can be converted back to intrinsic flux by inverting the operations. Bmaj and Bmin follow the same normalization scheme, with Bmaj clipped in the $[0.9,60.0]$ arcsec range and Bmin clipped in the $[0.3,30.0]$ arcsec range. They then go through a log10 function before being rescaled in the 0 to 1 range. For PA, it is crucial to consider whether the source is resolved. Trying to predict the position angle of an unresolved source would result in training noise. We first search for all sources with $\textrm{Bmaj}\leq 1.8$ arcsec and attribute them a target of $\hat{\textrm{PA}}=0$. This filtering only applies when defining the target, while nothing prevents the detector from predicting an angle for smaller sources at prediction time. We define two predicted parameters, $\cos{(\textrm{PA})}$ and $\sin{(\textrm{PA})}$. Considering that the target PA is in the $[-90,90]$ degree range, we linearly rescale $\sin{(\textrm{PA})}$ to the 0 to 1 range. Due to angular symmetries and degeneracies of PA, we obtained better results by predicting both $\cos{(\textrm{PA})}$ and $\sin{(\textrm{PA})}$ as output extra parameters, then reconstructing the predicted PA, than with a direct angle prediction. In summary, each box predicts the following vector for each source $x,y,w,h,P,O,f_{a},\textrm{Bmaj},\textrm{Bmin},\cos{(\textrm{PA})},\sin{(\textrm{PA})}\right>$. As discussed in Sect. 2.4, we can set independent $\gamma^{p}$ scaling values for all parameters to balance their respective importance in the loss. We list the $\gamma^{p}$ we used for all parameters in Table 3.
3.7.4 Remaining hyperparameters
For reproducibility purposes, we list all the hyperparameters for our application, which include those described in the appendix Appendix A.3 with the advanced description of the association process of YOLO-CIANNA. Understanding the specific behavior of each parameter should not be necessary to interpret the obtained result, and we refer to the Appendix when necessary.
@@ 3.7.4 Remaining hyperparameters
First, we use the ${\rm DIoU}$ as our matching metric (Appendix A.4), which implies that the $L^{{\rm fIoU}}$ limits can take values in the -1 to 1 range. Following our nine detection units definition with three scales, we set $\lambda_{\textrm{void}}=0.15$ for the six small detection units and $\lambda_{\textrm{void}}=0.01$ for the three larger ones. The $\lambda$ scaling factors, along with the preactivation scaling and limit values, are given in Table 2. The various $L^{{\rm fIoU}}$ and some other remaining hyperparameters are given in Table 3. For a more exhaustive view of all the detector hyperparameters, we recommend reading the provided example SDC1 scripts archived with CIANNA V-1.0.
Table 2: Scaling factors and limits related to the loss subparts.
@@
Table 3: Detection layer configuration hyperparameters.
Parameter $\gamma^{f_{a}}$ $\gamma^{\textrm{Bmaj}}$ $\gamma^{\textrm{Bmin}}$ $\gamma^{\cos{(\textrm{PA})}}$ $\gamma^{\sin{(\textrm{PA})}}$
@@
Parameter $L^{{\rm fIoU}}_{\textrm{GBNB}}$ $L^{{\rm fIoU}}_{\rm low}$ $L^{{\rm fIoU}}_{P}$ $L^{{\rm fIoU}}_{O}$ $L^{{\rm fIoU}}_{p}$
@@
Parameter $S_{ar}$ $\alpha_{\rm small}$ $N_{\rm rand}$ $\alpha_{\rm best}$ $\alpha_{\rm rand}$
@@
Table 4: First NMS rejection limit pairs.
$L^{{\rm fIoU}}_{\rm NMS}$ 0.05 -0.1 -0.3 -0.5
$L^{\rm obj}_{\rm NMS}$ 1.0 0.7 0.5 0.3
3.8 Network training
Figure 13: Evolution of the validation loss subparts during training, using the “natural” loss representation (Appendix A.7). The first 20 iterations are skipped. Optimized scores obtained using the procedure described in Sect. 3.9 are displayed every 200 iterations. The small difference in exact score that can be observed compared to Table 5 is due to rounding of the objectness thresholds, some minor adjustment of the largest prior threshold, and the use of FP16 mixed-precision at inference.
@@ 3.8 Network training
The layer weights are initialized to random values following a Glorot normal initialization (Glorot & Bengio, 2010), the batch size is set to $b_{s}=16$, and the numerical resolution is set to full 32-bit floating point for all network elements. The training images are dynamically generated and augmented from the training area, as described later in the current section. The learning rate starts low at around $3\times 10^{-6}$ and is increased linearly over the first 64000 training images up to the default value or $1.5{\times}10^{-4}$. Then, it follows an exponential decay of $3.13{\times}10^{-7}$ as a function of the number of training images up to a minimum value of $5\%$ of the default learning rate. A constant weight update momentum of 0.8 is also used. No weight decay is used since it always resulted in lower scores. The initial random association is set to $N_{\rm rand}=16000$ examples (Appendix A.6.2).
With our fully convolutional architecture, we could technically use any input size that is a multiple of the network reduction factor (see Appendix A.1). A too-small input size can cause many cutout edge effects, and a too-large input size reduces the number of independent examples we can generate from a finite training area, increasing the chances of overtraining. We observed that a training input size of $256{\times}256$ pixels resulted in the best achievable score for our specific application.
@@
To generate a training image, we draw a random pixel position inside the training area, extract a small region corresponding to our network input size, and apply pixel-preserving operations selected randomly from vertical or horizontal flips and $90$ or $-90$ degree rotations. Images from the additional noise region are selected at a low $5\%$ rate and follow the same augmentation rules. With this dynamic augmentation, it becomes impossible to define an epoch traditionally. By cutting out patches of the larger image, sources on the edges of a patch might be cut, which can negatively impact the detector capabilities (Sect. 3.4). We tried various approaches for handling these sources (Appendix A.6.4) but obtained the best scores with a simple exclusion of any target box that is not fully contained in the input patch. With this dynamic generation scheme, the classical epoch definition cannot be used. Therefore, to ease the monitoring of the training, we consider 1600 generated images to form a group that we call an iteration and use this quantity instead of the number of seen examples to express the advance of the network training.
@@
Because the labeled area is small, splitting it into independent training and validation subsets would significantly reduce the context diversity for both subsets. The detection performance would be negatively affected, and we would have no guarantee that the validation set is representative. For this reason, our training examples are randomly drawn from the full labeled area. Our validation dataset is built from the same region and catalog but using a fixed grid of nonoverlapping patches for a total of 100 images. We acknowledge that this validation set would fail to identify overtraining. However, we can rely on the SDC1 scorer to provide an independent metric on a separate dataset to evaluate our detector performances at a regular iteration interval.
@@
With this setup, the detector reaches a high score after only 400 iterations, but the extended sources are not properly constrained yet, and the global characterization is suboptimal. The detector usually requires around 3000 iterations to converge to its best score. After that, it oscillates for a few hundred iterations and eventually exhibits overtraining. We present the validation loss with scoring stamps every 200 iterations in Fig. 13. We note that using stricter values for parameters that control the association process like $L^{{\rm fIoU}}_{\textrm{GBNB}}$ or $L^{{\rm fIoU}}_{\rm low}$ would speed up the training. It would result in more stable results, but it also decreases the best achievable score by up to $2\%$.
@@
On an RTX 4090 GPU and by doing image augmentation on the CPU in parallel, we can reach almost 400 images per second of training performances. Reaching iteration 3000 takes around 3.5 hours. With the current training setup, the GPU memory footprint of the network is around 8GB. A save file representing the model at a given iteration requires 50MB of storage. We tried using mixed-precision training to leverage tensor-cores acceleration, but it always resulted in a score decrease of about $2\%$. Still, it speeds up network training by a factor of up to $1.8$ while preserving the relative performance impact of various hyperparameters, allowing a more efficient exploration of network architectures and hyperparameter combinations.
@@ 3.9 Prediction pipeline
To use our trained detector on the full SDC1 image, we need to decompose it into patches. As stated before, our method allows us to use any input size for the prediction, even if it differs from the one used for training. By decomposing the image into patches, some sources will end up close to the edges of a patch, resulting in poor detection or characterization. To overcome this issue at prediction time, we use overlapping patches with an overlap size of 32 pixels. This value was chosen to be half the maximum size for our target boxes. It is also a multiple of the reduction factor, making it an integer number of complete grid cells. With this setup, each source will likely be well-represented by at least one patch. In the case of multiple detections of the same source by two overlapping patches, having perfectly overlapping grid cells ensures that objectness scores are comparable. Thus, these multiple detections can be filtered through an inter-patch NMS process. From there, we observed that larger patches produce better results by reducing the proportion of sources close to an edge. We settled for a prediction input size of $512{\times}512$ pixels, which results in a good balance between prediction quality and compute efficiency. By setting a small offset on the edges of the full image, we can obtain a map composed of $67{\times}67$ partially overlapping patches.
@@
For each prediction patch, we first apply a per-detection-unit objectness threshold filtering. The threshold values are obtained through score optimization by identifying the objectness interval contributing positively to the score (Sect. 3.3 and Eq. 22). This optimization is only a refinement, and we would only lower our best score by a few percent by using naive guesses on threshold values instead of searching for the best ones. We also add a rejection criterion based on the average pixel flux inside the predicted box area as a function of the objectness score to exclude small false detections induced by very extended and bright sources. Multiple detections are then filtered by the NMS process for each patch (Sect. 2.8), using multiple $L^{{\rm fIoU}}_{\rm NMS}$ and $L^{\rm obj}_{\rm NMS}$ thresholds pairs that are given in Table 4. Following the NMS process, considering the best current score box, any other box that respects one of the condition pairs ${\rm DIoU}>L^{{\rm fIoU}}_{\rm NMS}$ and $O<L^{\rm obj}_{\rm NMS}$ is removed. With this setup, the rejection chance is based on a combination of the detection confidence and the distance to other confident detections. Multiple detections from overlapping patches are filtered using a secondary inter-patch NMS process. It works as the first NMS but with a constant $L^{{\rm fIoU}}_{\rm NMS}=-0.15$. From the filtered box list, we construct a catalog in the SDC1 scorer format by inverting our normalizations on the predicted parameters.
@@
On an RTX 4090 GPU, the raw prediction compute performance at the largest available batch size is around 300 images per second for an input size of $512{\times}512$ pixels, or almost 80 million pixels per second. The raw processing time for the full 4GB image of 32768 square pixels after the network loading is around 15 seconds. To this, we need to add the post-processing time of the prediction pipeline, which strongly depends on the minimum objectness threshold values for the different detection units. Using an end-to-end Python pipeline that produces the converted detection catalog from the raw detection catalog, the post-processing time is about 40s on a single CPU core (Ryzen 9 5900X). Interestingly, the detection performances are fully preserved when doing the prediction in mixed-precision using 16-bit numerical computations with a 32-bit accumulator when necessary, allowing the use of Nvidia AI-dedicated tensor-cores compute units available on modern GPU architectures. With this, we can reach around 500 images of $512{\times}512$ pixels per second, or around 130 million pixels per second, of raw computing performance. At this point, the time required to save the results on our high-end SSD storage becomes the limiting factor, and the network loading time becomes comparable to the prediction time over the full image. This confirms that our method is suited for application over large surveys and real-time detection (Sect. 5.3).